a block of mass5kg is resting on a smooth surface.at what angle a force of 20N be acted on the body so that it will acqurie a kinetic energy of 40J AFTER MOVING 4M?

Grade:12th pass

1 Answers

Sujit Kumar
6 years ago
I’m not sure if my answer is correct. According to me..... mass of the block is not required..... unless you want to calculate the acceleration.......... but this is how i did it....
 
Energy = Force x Distance
Hence the force that acted on the block = 40/4 = 10N
 
Now consider the block is being pulled by a thread with an angle \Theta with the horizontal.
 
 
therefore
 Cos\Theta=\frac{Adjacent Side}{Hypotenuse}=\frac{10}{20}=\frac{1}{2}
 
=>Cos\Theta=\frac{1}{2}
 =>\Theta=60^{\circ}
 
Ans: The force of 20N should be acted on the object with an angle of 60°